OpenStudy (butterflyhope):

A budget calls for $200 less than one-third of the clothing portion of the budget to be spent on the man of the house. If the man's needs are estimated at $600 for the year, how much should be budgeted for the entire family for the year? Thanks in advance :)

OpenStudy (anonymous):

y=1/3x-200

OpenStudy (anonymous):

600=1/3x-200

OpenStudy (anonymous):

800=1/3x

OpenStudy (anonymous):

3/1 * 800 = 1/3x * 3/1

OpenStudy (anonymous):

2400=x

OpenStudy (anonymous):

$2400 should be budgeted for the entire family for the year.

OpenStudy (anonymous):

Do you understand how we got that answer?

OpenStudy (anonymous):

To check, you'd plug 2400 into the original problem. 600=1/3(2400)-200

OpenStudy (anonymous):

600=800-200

OpenStudy (anonymous):

600=600

OpenStudy (butterflyhope):

I understand, thank you :D